If a bike is going 24 km/h and a car is going 60 km/h but the car is 70km behind how long would it take for the car to catch up
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

If a bike is going 24 km/h and a car is going 60 km/h but the car is 70km behind
How long would it take for the car to catch up?
---------------
Bike DATA:
rate = 24 km/hr ; distance = x km ; time = d/r = x/24 hrs
-------
Car DATA:
rate = 60 km/hr ; distance = x+70 km ; time = (x+70)/60 hrs
------
Equation:
time = time
x/24 = (x+70)/60
60x = 24x + 24*70
36x = 24*70
x = 140/3
---
Ans: time = (140/3)/24 = 140/72 = 1.94 hrs = 1 hr 56 min 24 seconds
------------------
Cheers,
Stan H.
================

Answer by richwmiller(17219)

24x+70=60x
70=36x
70/36=x
x=1.94 hours which is 1 hr 56 min 24 seconds